# The equivalence of asymptotic average shadowing and vague specification properties and its consequences

### v1

@@ Abstract
We establish the equivalence between the asymptotic average shadowing and the vague specification properties and we use this equivalence to answer a question posed by Downarowicz and Weiss [Ergod. Th. & Dynam. Sys. 44(9) (2024), 2565-2580]. Additionally, we prove that the asymptotic average shadowing property is equivalent to the average shadowing property if the phase space is complete with respect to the dynamical Besicovitch pseudometric. We use the latter result to prove that proximal and minimal shift spaces from [Minimal and proximal examples of $\bar{d}$-stable and $\bar{d}$-approachable shift spaces, Ergod. Th. & Dynam. Sys., (2024) First View] possess the asymptotic average shadowing property.
@@ 1 Introduction
00footnotetext: Mathematics Subject Classification 2020. Primary 37B65 ; Secondary 37B05, 37B1000footnotetext: Key words. shadowing property, specification property, Besicovitch pseudometric, shift space.
The shadowing property, also called the pseudo-orbit tracing property, is a fundamental concept in the theory of dynamical systems. Informally, it means that every pseudo-orbit (a sequence of points in the phase space such that every point is close to the image of the previous one) is traced by an actual orbit. Such property played a crucial role in the development of the theory of hyperbolic dynamical systems because of the Shadowing Lemma which is a key step in the proof of structural stability of hyperbolic sets (see [4, 12]).
Besides its applications to differentiable dynamics, the shadowing property became a subject of studies in topological dynamics. This led to an introduction of different notions of pseudo-orbits and tracing, resulting in variants of the shadowing property, such as the limit shadowing property [8], the average shadowing property [3], and the asymptotic average shadowing property [9].
@@
In the early 70’s Bowen [4] introduced another notion of tracing of sequence of points in the space by an orbit. Bowen called it the specification property, and used it to study Axiom $A$ diffeomorphisms. Roughly speaking, the specification property guarantees the existence of an orbit to trace arbitrary, but finite, collection of segments of orbits provided that these segments are sparse enough in time. Systems with the specification property have very rich dynamics and the property has been used to investigate other properties like intrinsic ergodicity, topological mixing, positive topological entropy and others (see [5, 20]).
In order to study dynamical systems without Bowen’s specification, weaker notions were introduced, such as almost specification, weak specification and relative specification (we refer [16] for an overview).
The initial goal of this study was to answer a question from a recent work of Downarowicz and Weiss [7]. They asked about the relation between the weak specification property and a less explored variant of Bowen’s specification, called the vague specification property. The latter was introduced by Kamae [11]. We prove that the weak specification property implies the vague specification property and the converse does not hold. The conclusion follows from our main result, that establishes the equivalence of the asymptotic average shadowing and the vague specification properties, together with known relations between the weak specification and the asymptotic average shadowing properties.
@@
More precisely, in [17], it is shown that both almost specification and weak specification properties imply the asymptotic average shadowing property. It is also known that there are systems with almost specification but without weak specification and vice-versa (see [14, 18]). As a consequence of the equivalence of the vague specification and the asymptotic average shadowing properties, one sees that the weak specification property implies the vague specification property and that the systems with the almost specification property but without the weak specification property serve as counterexamples for the converse. Therefore, the vague specification property can be considered one of the weakest forms of specification (see [16, 17]).
The equivalence of the vague specification and the asymptotic average shadowing properties also allows us to answer questions about the latter property. Indeed, we can use Kamae’s result (Proposition 5 in [11]), which shows that the vague specification property is inherited by factors, to address Question 10.6 in [14], confirming that the asymptotic average shadowing property is inherited by factors.
@@
When studying the asymptotic average shadowing property, it is natural to ask its relation with the average shadowing property. It is already known that the asymptotic average shadowing property implies the average shadowing property. This result was first proved for surjective topological dynamical systems by Kulczycki, Kwietniak and Oprocha in [14]. Later, the result was improved by Wu, Oprocha and Chen in [22], who showed that the assumption of surjectivity could be dropped.
@@
Roughly speaking, the dynamical Besicovitch pseudometric, defined on the phase space of a topological dynamical system, measures the asymptotic average distance between the iterates of two initial points. We show that for a surjective topological dynamical system $(X,T)$, the asymptotic average shadowing property is equivalent to $(X,T)$ being complete with respect to the dynamical Besicovitch pseudometric and having the average shadowing property. This provides a step toward answering Question 10.3 in [14].
Despite our efforts, we were unable to find any examples of systems that have the average shadowing property but are not Besicovitch complete. Thus, Question 10.3 in [14] remains unsolved and can be restated as: “Does the average shadowing property imply Besicovitch completeness?”
For surjective shift spaces, the average shadowing property is equivalent to the $\bar{d}$-shadowing property, introduced by Konieczny, Kupsa and Kwietniak in [13]. This equivalence was claimed without a proof in [13]. Since we rely on this equivalence in our work, we provide a proof of this statement for the sake of completeness.
@@
We recall that the $\bar{d}$ and the Besicovitch pseudometrics are uniformly equivalent on shift spaces (see [17]). As a consequence, we obtain that, for surjective shift spaces, if we assume completeness with respect to the $\bar{d}$ pseudometric, then the $\bar{d}$-shadowing and the asymptotic average shadowing properties are equivalent.
@@
In [6], the authors constructed a proximal shift space and a class of minimal shift spaces with the $\bar{d}$-shadowing property. The authors acknowledged Piotr Oprocha for the ideas that led to the construction of the minimal shift spaces. We prove that these examples are $\bar{d}$-complete, and therefore, as they are surjective, they have the asymptotic average shadowing property. This allows us to answer Question 10.2 in [14], by presenting minimal systems with the asymptotic average shadowing property.
@@
Specification-like properties are the most common tools used to show that ergodic measures are entropy-dense in the space of invariant measures (see [16, 19, 20] for definitions and details). The proximal example and the class of minimal examples constructed in [6] are shown to have ergodic measures entropy-dense in the space of invariant measures. The surprise about such systems was that they do not seem to manifest any specification-like property, and therefore, different strategies are used to obtain the result. However, our results confirm that all these examples possess the vague specification property, which suggests that the vague specification property may imply entropy-density of ergodic measures.
@@ 2 Preliminaries
We write =$\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For every $A\subset\mathbb{N}_{0}$, we denote by $|A|\in\mathbb{N}_{0}\cup\{\infty\}$ the cardinality of $A$. For $n,m\in\mathbb{N}_{0}$ we write $[n,m)=\{i\in\mathbb{N}_{0}:n\leq i<m\}$, $(n,m]=\{i\in\mathbb{N}_{0}:n<i\leq m\}$ and $[n,m]=\{i\in\mathbb{N}_{0}:n\leq i\leq m\}$.
@@
We denote by $X^{\infty}$ the family of all $X$-valued sequences indexed by $\mathbb{N}_{0}$, that means,
@@
Typically, we write $\underline{x}=\{x_{n}\}_{n=0}^{\infty}$ for elements of $X^{\infty}$. We denote by $S$ the shift operator acting on $X^{\infty}$, that is, $S(\underline{x})=\{x_{n+1}\}_{n=0}^{\infty}$. We always endow $X^{\infty}$ with the product topology, which is compatible with the metric
@@
A pair $(X,T)$ is said to be a topological dynamical system if $T\colon X\to X$ is a continuous map. The orbit of a point $x\in X$ under $T$ is usually defined to be the set $\{T^{n}(x)\in X:n\in\mathbb{N}_{0}\}$ but, for our purposes, it is more convenient to treat orbits as sequences and we write $\underline{x}_{T}=\{T^{n}(x)\}_{n=0}^{\infty}\in X^{\infty}$. We write $\Gamma_{T}=\{\underline{x}_{T}\in X^{\infty}:x\in X\}$ for the space of all orbits under $T$. It is not hard to see that $\Gamma_{T}$ is a closed subspace of $X^{\infty}$, and therefore, $\Gamma_{T}$ is compact.
@@
It is easy to see that the pseudometrics $\rho_{B}$ and $\rho_{B}^{\prime}$ on $X^{\infty}$ are uniformly equivalent (see Lemma 2 in [17] for a proof).
@@
Let $(X,T)$ be a topological dynamical system. We say that a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ is Besicovitch-Cauchy if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every $n,m\geq N$ it holds
@@
The notion of Cauchy sequences and completeness for the Besicovitch pseudometric on $X^{\infty}$ could also be defined accordingly, and replicating the proof of Proposition 2 in [2] we see that $X^{\infty}$ is always Besicovitch complete.
3 Pseudo-orbits for $(X,T)$
@@
A sequence $\underline{x}\in X^{\infty}$ is called a $\delta$-pseudo-orbit for $(X,T)$ if $\rho(T(x_{n}),x_{n+1})<\delta$ for every $n\in\mathbb{N}_{0}.$
Allowing the errors to be small in different sense than the uniform one, like vanishing in the limit or on average, motivated definitions of different types of pseudo-orbits (we refer [9, 14, 19, 16] for more details). One of these generalized notions is the asymptotic pseudo-orbit.
@@
A sequence $\underline{x}\in X^{\infty}$ is called an asymptotic pseudo-orbit for $(X,T)$ if
@@
Blank [3] analyzed pseudo-orbits, called average pseudo-orbits, in which the average of the errors becomes small in a long term.
@@
A sequence $\underline{x}\in X^{\infty}$ is said to be a $\delta$-average pseudo-orbit for $(X,T)$ if there exists $N\in\mathbb{N}$ such that for every $n\geq N$ and $k\in\mathbb{N}_{0}$ it holds
@@
From Definitions 3.2 and 3.3 it is not hard to see that asymptotic pseudo-orbits are $\delta$-average pseudo-orbits for any arbitrary $\delta>0$.
@@
If $\underline{x}\in X^{\infty}$ is an asymptotic pseudo-orbit for $(X,T)$, then $\underline{x}$ is a $\delta$-average pseudo-orbit for $(X,T)$ for every $\delta>0$.
Later, Gu [9] studied the case when the average of the errors vanishes in the limit.
@@
A sequence $\underline{x}\in X^{\infty}$ is said to be an asymptotic average pseudo-orbit for $(X,T)$ if
@@
A sequence $\underline{x}\in X^{\infty}$ is an asymptotic average pseudo-orbit for $(X,T)$ if and only if for every $\varepsilon>0$ it holds
@@
The combination of Lemmas 3.1 and 3.3 in [16] implies the following proposition.
Proposition 3.7.
Let $(X,T)$ be a surjective topological dynamical system. For every asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$ there exists an asymptotic pseudo-orbit $\underline{z}\in X^{\infty}$ such that
@@
In order to define the vague specification property, Kamae [11] studied an unnamed family of sequences in $X^{\infty}$ that we will call vague pseudo-orbit.
@@
We call $\underline{x}\in X^{\infty}$ a vague pseudo-orbit for $(X,T)$ if for any open neighborhood ${\mathcal{U}}$ of $\Gamma_{T}$ in $X^{\infty}$ it holds
@@
It turns out that asymptotic average pseudo-orbits and vague pseudo-orbits are the same sequences. In order to prove the equivalence between asymptotic average pseudo-orbits and vague pseudo-orbits we will need some lemmas.
@@
Let $G\subset\mathbb{N}_{0}$. If $d(G)=1$, then for every $k\in\mathbb{N}$ it holds
@@
Let $G\subset\mathbb{N}_{0}$ be such that $d(G)=1$ and $k\in\mathbb{N}$. Set $\tilde{J}=\mathbb{N}_{0}\backslash G$. Note that $d(G)=1$ implies $d(\tilde{J})=0$, that is,
$\lim_{n\to\infty}\frac{|\tilde{J}\cap[0,n)|}{n}=0.$
Set
$J=\bigcup_{j\in\tilde{J}}[j-k,j].$
By construction, we have $|J\cap[0,n)|\leq k|\tilde{J}\cap[0,n)|+k$ for every $n\in\mathbb{N}$. Consequently,
$d(J)=\lim_{n\to\infty}\frac{|J\cap[0,n)|}{n}\leq\lim_{n\to\infty}\frac{k|\tilde{J}\cap[0,n)|+k}{n}=0.$
Lastly, observe that
$\{n\in\mathbb{N}_{0}:[n,n+k)\subset G\}=\mathbb{N}_{0}\backslash J.$
Therefore,
$d(\{n\in\mathbb{N}_{0}:[n,n+k)\subset G\})=1.\qed$
Straightforward from the definition of $\rho^{\infty}$ we have the following lemma.
@@
By compactness of $\Gamma_{T}$, there exist $\varepsilon>0$ and $x^{(1)},\ldots,x^{(m)}\in X$ such that
$\Gamma_{T}\subset\bigcup_{i=1}^{m}B^{\infty}(x^{(i)},\varepsilon)\subset\bigcup_{i=1}^{m}B^{\infty}(x^{(i)},2\varepsilon)\subset{\mathcal{U}}.$
Let $k\in\mathbb{N}$ be such that $\frac{1}{2^{k}}<\varepsilon$. Take $\underline{z}\in X^{\infty}$ so that
@@
By Lemma 3.10, we have $\rho^{\infty}(\underline{z},\underline{z_{0}}_{T})<\varepsilon.$ Note that $\underline{z_{0}}_{T}\in B^{\infty}(x^{(i)},\varepsilon)$ for some $i\in[1,m]$. Therefore, by the triangle inequality, $\underline{z}\in B^{\infty}(x^{(i)},2\varepsilon)\subset{\mathcal{U}}$. ∎

### v2

# On the weakness of the vague specification property

@@ Abstract
We show that the vague specification property is strictly weaker than most of the specification-like properties, by establishing its equivalence with the asymptotic average shadowing property. In particular, we see that the weak specification property implies the vague specification property, but the converse does not hold, answering the question posed by Downarowicz and Weiss in [Ergod. Th. & Dynam. Sys. 44(9) (2024), 2565-2580]. Additionally, we prove that, for surjective systems, the asymptotic average shadowing property is equivalent to the average shadowing property if the phase space is complete with respect to the dynamical Besicovitch pseudometric. We use the combination of both results to prove that the proximal and minimal shift spaces from [Ergod. Th. & Dynam. Sys., 45(2) (2025), 396–426] possess the vague specification property (asymptotic average shadowing property). Our findings also allow us to address a couple of questions from [Fund. Math., 224(3) (2014), 241–278] about the asymptotic average shadowing property.
@@ 1 Introduction
00footnotetext: Mathematics Subject Classification 2020. Primary 37B65 ; Secondary 37B05, 37B1000footnotetext: Key words. shadowing property, specification property, Besicovitch pseudometric, shift space, minimal system, proximal system.00footnotetext: MEC was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885 and NCN Polonez Bis grant No. 2022/47/P/ST1/00854. AT was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885 and NCN Sonata Bis grant no. 2019/34/E/ST1/00082.
@@
In the early 1970s, Bowen [Bow71] introduced the specification property and used it to study Axiom A diffeomorphisms. Roughly speaking, the specification property guarantees the existence of an orbit to trace arbitrary, but finite, collections of segments of orbits, provided that these segments are sparse enough in time. Systems with the specification property have very rich dynamics since this property implies, for example, topological mixing, positive topological entropy, and intrinsic ergodicity (see [Bow74, Sig74]).
Although the specification property is satisfied by important classes of systems, such as mixing interval maps and mixing shifts of finite type, weaker forms of specification were introduced in order to cover relevant classes without specification in the sense of Bowen. Some examples of these variants are the weak specification property [Mar80, Dat83], satisfied by all automorphisms of compact Abelian groups for which the Haar measure is ergodic [Dat90] and the almost specification [PS05, Tho12], satisfied by $\beta$-shifts. Specification-like properties have been studied by several authors, and we refer to [KŁO16] for an overview. Below, we present a simplified version of the diagram from [KŁO16] and emphasize that there are no additional implications between these notions besides those given by transitivity.
specificationperiodic specificationalmostspecificationweak specificationperiodicweak specification
The specification property and its variants have been extensively used in ergodic theory to investigate, for example, measures of maximal entropy, large deviation, and the existence of generic points for not necessarily ergodic measures. Recently, Downarowicz and Weiss [DW24] used the weak specification to prove results about lifting generic points. Kamae [Kam76] obtained similar results in 1976 under the assumption of a less explored specification-like property (not considered in [KŁO16]), called the vague specification property, which he introduced in the same work. This connection led to the question posed at the end of Downarowicz and Weiss’ work about how the weak and the vague specification properties relate to each other.
We show that the vague specification property is implied by all the properties displayed in the diagram above, providing, as a consequence, an answer to Downarowicz and Weiss’ question. This follows from Theorem A, where we establish the equivalence of the vague specification and a variant of the shadowing property introduced by Gu [Gu07], called the asymptotic average shadowing property, combined with previous results on the latter property.
Theorem A.
@@
In [KŁO17], it is shown that both the almost specification and the weak specification properties imply the asymptotic average shadowing property. Therefore, we conclude that the vague specification property is one of the weakest forms of specification (see [KŁO16, KŁO17]) and we can complete the diagram above by adding the vague specification property in the following way.
specificationperiodic specificationalmostspecificationweak specificationperiodicweak specificationvaguespecification
Furthermore, while both the weak and the almost specification properties imply the existence of several pairwise disjoint invariant sets, we show that the vague specification property can be satisfied by proximal and minimal systems, which reinforces its weakness. Indeed, we prove that the proximal example and the class of minimal shift spaces constructed in [CKKK24] possess the vague specification property.
Theorem B.
There are non-trivial minimal and non-trivial proximal systems which have the vague specification property.
Theorems A and B originated from our studies on variants of the classical shadowing property. The shadowing property, also known as pseudo-orbit tracing property, played a crucial role in the development of the theory of hyperbolic dynamical systems due to the Shadowing Lemma, which is a key step in the proof of structural stability of hyperbolic sets (see [Bow71, KH95]). Informally, the shadowing property, means that every pseudo-orbit (a sequence of points in the phase space such that every point is close to the image of the previous one) is traced by an actual orbit.
Besides its applications to differentiable dynamics, the shadowing property became a subject of studies in topological dynamics. This led to the introduction of different notions of pseudo-orbits and tracing, resulting in variants of the shadowing property, such as the limit shadowing property [ENP97], the average shadowing property [Bla88], and the asymptotic average shadowing property [Gu07].
While the original notion of pseudo-orbit consists of small errors at each iterate, in the asymptotic average shadowing property, the type of pseudo-orbits considered requires the average of these errors to vanish in the limit. Similarly, the notion of tracing is also defined by the average distance between the iterates of a point and the terms of the pseudo-orbits vanishing in the limit. The investigation of the connection between the asymptotic average shadowing and the vague specification properties, which led to Theorem A, was motivated by the fact that the concept of tracing considered in both properties, although differently stated, is actually the same. The notion of tracing used on the vague specification property is defined via the asymptotic density of natural numbers.
@@
The average shadowing property, introduced by Blank in [Bla88], and recently revisited by himself in [Bla24], was an inspiration for the asymptotic average shadowing property. Thus, when studying the asymptotic average shadowing property, it is natural to ask its relation with the average shadowing property. It is already known that the asymptotic average shadowing property implies the average shadowing property. This result was first proved for surjective topological dynamical systems by Kulczycki, Kwietniak, and Oprocha in [KKO14]. Later, the result was improved by Wu, Oprocha and Chen in [WOC16], who showed that the assumption of surjectivity could be dropped. In [KKO14], one of the questions posed by the authors ask whether the converse is true (Question 10.3 in [KKO14]).
@@
In Theorem C we provide a step toward answering Question 10.3 in [KKO14], by showing that for a surjective topological dynamical system $(X,T)$, the asymptotic average shadowing property is equivalent to $(X,T)$ being complete with respect to the dynamical Besicovitch pseudometric and having the average shadowing property. Roughly speaking, the dynamical Besicovitch pseudometric, defined on the phase space of a topological dynamical system, measures the asymptotic average distance between the iterates of two initial points. It is worth noting that some authors define the dynamical Besicovitch pseudometric without naming it, and we decided to adopt this terminology in accordance with [KŁO17].
Theorem C.
Although we believe that Theorem C is interesting on its own, we also stress that this result is crucial for the proof of Theorem B. In fact, we use an equivalent version of Theorem C for shift spaces, that is, subsystems of the full shift. Several dynamical properties, such as topological mixing and the specification property, have simpler formulations in the context of shift spaces. Konieczny, Kupsa, and Kwietniak [KKK23] introduced the $\bar{d}$-shadowing property, a variant of the shadowing property for shift spaces that is strongly related to the average shadowing property. Indeed, the average shadowing property implies the $\bar{d}$-shadowing property, and the converse holds for surjective shift spaces. These relations were claimed without a proof in [KKK23], as we rely on this equivalence in our work, for the sake of completeness, we present a proof for these statements.
Theorem D.
Let $X$ be a surjective shift space. Then $X$ has the average shadowing property if and only if $X$ has the $\bar{d}$-shadowing property.
@@
We recall that the $\bar{d}$ and the Besicovitch pseudometrics are uniformly equivalent on shift spaces (see [KŁO17]). As a consequence, Theorem C implies that, for surjective shift spaces, if we assume completeness with respect to the $\bar{d}$ pseudometric, then the $\bar{d}$-shadowing and the asymptotic average shadowing properties are equivalent. In particular, from Theorem A, we obtain the following corollary.
Corollary.
Let $X$ be a surjective shift space. Then $X$ has the vague specification property if and only if $X$ is $\bar{d}$-complete and has the $\bar{d}$-shadowing property.
@@
In [CKKK24], the authors constructed a proximal shift space and a class of minimal shift spaces with the $\bar{d}$-shadowing property. The authors acknowledged Piotr Oprocha for the ideas that led to the construction of the minimal shift spaces. We prove that these examples are $\bar{d}$-complete, and therefore, as they are surjective, they have the vague specification property which proves Theorem B.
@@
Specification-like properties are the most common tools used to prove entropy-density of ergodic measures in the space of invariant measures (see [KŁO16, Pav16, Sig74] for definitions and details). The proximal example and the class of minimal shift spaces constructed in [CKKK24] are shown to have ergodic measures entropy-dense in the space of invariant measures. The surprise about such systems was that they do not seem to manifest any specification-like property and, therefore, different strategies are used to obtain the result. However, our results confirm that all these examples possess the vague specification property, which suggests that the vague specification property may imply entropy-density of ergodic measures.
Lastly, we also use the equivalence of the vague specification and the asymptotic average shadowing properties to answer other questions about the latter property raised in [KKO14]. Indeed, the minimal shift space from Theorem B answers Question 10.2 in [KKO14], confirming that systems with the asymptotic average shadowing property may be minimal. Moreover, we can use Kamae’s result (Proposition 5 in [Kam76]), which shows that the vague specification property is inherited by factors, to address Question 10.6 in [KKO14], confirming that the asymptotic average shadowing property is inherited by factors.
Here is the organization of the paper. In Section 2, we present some basic notions and notation used throughout this work. In Section 3, we analyze different notions of pseudo-orbits and the relations between them, and we prove the equivalence between asymptotic average pseudo-orbits and vague pseudo-orbits, which is the key step to prove Theorem A. In Section 4, we prove Theorems A and C. In Section 5, we present the basic definitions of shift spaces, we study the $\bar{d}$ pseudometric and $\bar{d}$-shadowing property, and we prove Theorem D. Section 6 is devoted to the proof of Theorem B. Finally, in Section 7, we discuss some consequences of our results.
@@ 2 Preliminaries
We write =$\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $A\subseteq\mathbb{N}_{0}$ and $i\in\mathbb{N}_{0}$, we write $A-i=\{n\in\mathbb{N}_{0}:n+i\in A\}$. For every $A\subset\mathbb{N}_{0}$, we denote by $|A|\in\mathbb{N}_{0}\cup\{\infty\}$ the cardinality of $A$. For $n,m\in\mathbb{N}_{0}$ we write $[n,m)=\{i\in\mathbb{N}_{0}:n\leq i<m\}$, $(n,m]=\{i\in\mathbb{N}_{0}:n<i\leq m\}$ and $[n,m]=\{i\in\mathbb{N}_{0}:n\leq i\leq m\}$.
@@
We denote by $X^{\infty}$ the family of all $X$-valued sequences indexed by $\mathbb{N}_{0}$, that is,
@@
Typically, we write $\underline{x}=\{x_{n}\}_{n=0}^{\infty}$ for the elements of $X^{\infty}$. We denote by $S$ the shift operator acting on $X^{\infty}$, that is, $S(\underline{x})=\{x_{n+1}\}_{n=0}^{\infty}$. We always endow $X^{\infty}$ with the product topology, which is compatible with the metric
@@
A pair $(X,T)$ is said to be a topological dynamical system if $T\colon X\to X$ is a continuous map. The orbit of a point $x\in X$ under $T$ is usually defined as the set $\{T^{n}(x)\in X:n\in\mathbb{N}_{0}\}$ but, for our purposes, it is more convenient to treat orbits as sequences and we write $\underline{x}_{T}=\{T^{n}(x)\}_{n=0}^{\infty}\in X^{\infty}$. We write $\Gamma_{T}=\{\underline{x}_{T}\in X^{\infty}:x\in X\}$ for the space of all orbits under $T$. It is not hard to see that $\Gamma_{T}$ is a closed subspace of $X^{\infty}$, and therefore, $\Gamma_{T}$ is compact.
@@
It is easy to see that the pseudometrics $\rho_{B}$ and $\rho_{B}^{\prime}$ on $X^{\infty}$ are uniformly equivalent (see Lemma 2 in [KŁO17] for a proof).
@@
Proof.
Fix $\underline{x},\underline{z}\in X^{\infty}$. Assume that for every $\varepsilon>0$ we have
$d(\{n\in\mathbb{N}:\rho(x_{n},z_{n})<\varepsilon\})=1.$
Then by definition $\rho_{B}^{\prime}(\underline{x},\underline{z})=0$.
To prove the converse, assume that $\rho_{B}^{\prime}(\underline{x},\underline{z})=0$ and fix $\varepsilon>0$. We observe that for each $0<\delta<\varepsilon$ we have
$\{n\in\mathbb{N}:\rho(x_{n},z_{n})<\delta\}\subseteq\{n\in\mathbb{N}:\rho(x_{n},z_{n})<\varepsilon\}.$
Fix $0<\delta<\varepsilon$. Since $\rho_{B}^{\prime}(\underline{x},\underline{z})<\delta$, one has
$1-\delta\leq\underline{d}(\{n\in\mathbb{N}:\rho(x_{n},z_{n})<\delta\})\leq\underline{d}(\{n\in\mathbb{N}:\rho(x_{n},z_{n})<\varepsilon\}).$ (1)
Since (1) holds for every $0<\delta<\varepsilon$, we conclude that
$d(\{n\in\mathbb{N}:\rho(x_{n},z_{n})<\varepsilon\})=1.\qed$
@@
Let $(X,T)$ be a topological dynamical system. We say that a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ is Besicovitch-Cauchy if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every $n,m\geq N$ we have
@@
The notion of Cauchy sequences and completeness for the Besicovitch pseudometric on $X^{\infty}$ could also be defined accordingly, and replicating the proof of Proposition 2 in [BFK97] we see that $X^{\infty}$ is always Besicovitch complete.
3 Pseudo-orbits in topological dynamical systems
Usually, pseudo-orbits are defined together with their corresponding shadowing-like properties. In our case, we will analyze some pseudo-orbits on their own, apart from their corresponding tracing properties. To avoid unnecessary definitions, we decided to devote this section to generalizations of pseudo-orbits and their relations.
@@ 3 Pseudo-orbits in topological dynamical systems
A sequence $\underline{x}\in X^{\infty}$ is called a $\delta$-pseudo-orbit in $(X,T)$ if $\rho(T(x_{n}),x_{n+1})<\delta$ for every $n\in\mathbb{N}_{0}.$
Allowing the errors to be small in a different sense than the uniform one, such as vanishing in the limit or on average, lead to definitions of different types of pseudo-orbits (we refer to [Bla88, ENP97, Pil99, Gu07] for more details). One of these generalized notions is the asymptotic pseudo-orbit.
@@
A sequence $\underline{x}\in X^{\infty}$ is called an asymptotic pseudo-orbit in $(X,T)$ if
@@
Blank [Bla88] analyzed pseudo-orbits, nowadays called average pseudo-orbits, in which the average of the errors becomes small in the long term.
@@
A sequence $\underline{x}\in X^{\infty}$ is said to be a $\delta$-average pseudo-orbit in $(X,T)$ if there exists $N\in\mathbb{N}$ such that for every $n\geq N$ and $k\in\mathbb{N}_{0}$ it holds that
@@
From Definitions 3.2 and 3.3 it is not hard to see that asymptotic pseudo-orbits are $\delta$-average pseudo-orbits for any $\delta>0$.
@@
If $\underline{x}\in X^{\infty}$ is an asymptotic pseudo-orbit in $(X,T)$, then $\underline{x}$ is a $\delta$-average pseudo-orbit in $(X,T)$ for every $\delta>0$.
Later, Gu [Gu07] studied the case where the average of errors vanishes in the limit.
@@
A sequence $\underline{x}\in X^{\infty}$ is said to be an asymptotic average pseudo-orbit in $(X,T)$ if
@@
A sequence $\underline{x}\in X^{\infty}$ is an asymptotic average pseudo-orbit in $(X,T)$ if and only if for every $\varepsilon>0$ the following holds
@@
Asymptotic pseudo-orbits and asymptotic average pseudo-orbits are different. However, if a topological dynamical system is chain mixing, then given an asymptotic average pseudo-orbit one can obtain an asymptotic pseudo-orbit by making small modifications.
Definition 3.7.
Proposition 3.8 (Lemma 3.3 in [KKO14]).
Let $(X,T)$ be a chain mixing topological dynamical system. For every asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$ there exists an asymptotic pseudo-orbit $\underline{z}\in X^{\infty}$ such that
@@
In order to define the vague specification property, Kamae [Kam76] studied an unnamed family of sequences in $X^{\infty}$ which we will call vague pseudo-orbit.
@@
We call $\underline{x}\in X^{\infty}$ a vague pseudo-orbit in $(X,T)$ if for any open neighborhood ${\mathcal{U}}$ of $\Gamma_{T}$ in $X^{\infty}$ we have
@@
It turns out that asymptotic average pseudo-orbits and vague pseudo-orbits are the same sequences. In order to prove the equality between asymptotic average pseudo-orbits and vague pseudo-orbits we will need some lemmas.
@@
Let $G\subset\mathbb{N}_{0}$. If $d(G)=1$, then for every $k\in\mathbb{N}$ the following holds
@@
Let $G\subset\mathbb{N}_{0}$ be such that $d(G)=1$ and $k\in\mathbb{N}$. Set $G_{k}=\{n\in\mathbb{N}_{0}:[n,n+k)\subset G\}$. It is enough to prove that $\bar{d}(\mathbb{N}_{0}\setminus G_{k})=0$.
Note that $G_{k}=\bigcap_{i=0}^{k-1}(G-i).$ Note also that $d(G)=d(G-i)$ for every $i\in[0,k)$. As a consequence, we have $\bar{d}(\mathbb{N}_{0}\setminus(G-i))=0$ for every $i\in[0,k)$. Lastly, since $\bar{d}(A\cup B)\leq\bar{d}(A)+\bar{d}(B)$ for every $A,B\subseteq\mathbb{N}_{0}$, we complete the proof by observing that
$\mathbb{N}_{0}\setminus G_{k}=\bigcup_{i=0}^{k-1}\Big{(}\mathbb{N}_{0}\setminus\left(G-i\right)\Big{)},$
which implies, $\bar{d}(\mathbb{N}_{0}\setminus G_{k})=0$. ∎
Directly from the definition of $\rho^{\infty}$ we have the following lemma.
@@
Fix ${\mathcal{U}}\subset X^{\infty}$ an open set containing $\Gamma_{T}$. For each $x\in X$ there exists $\varepsilon=\varepsilon(x)>0$ such that $B^{\infty}(x,2\varepsilon)\subset{\mathcal{U}}$. By compactness of $\Gamma_{T}$, there exist $x^{(1)},\ldots,x^{(m)}\in X$ and $\{\varepsilon_{1},\ldots,\varepsilon_{m}\}$ such that
$\Gamma_{T}\subset\bigcup_{i=1}^{m}B^{\infty}(x^{(i)},\varepsilon_{i})\subset\bigcup_{i=1}^{m}B^{\infty}(x^{(i)},2\varepsilon_{i})\subset{\mathcal{U}}.$
Let $\varepsilon=\min\{\varepsilon_{1},\ldots,\varepsilon_{m}\}$ and $k\in\mathbb{N}$ be such that $\frac{1}{2^{k}}<\varepsilon$. Take $\underline{z}\in X^{\infty}$ such that
@@
By Lemma 3.11, we have $\rho^{\infty}(\underline{z},\underline{z_{0}}_{T})<\varepsilon.$ Note that $\underline{z_{0}}_{T}\in B^{\infty}(x^{(i)},\varepsilon_{i})$ for some $i\in\{1,\ldots,n\}$. Therefore, by the triangle inequality, $\underline{z}\in B^{\infty}(x^{(i)},2\varepsilon_{i})\subset{\mathcal{U}}$. ∎